OpenStudy (anonymous):

How to solve -40=m-10m

OpenStudy (anonymous):

According to your question m-10m=-40 -9m=-40 so we can write 9m=40 ie m=40/9 m=4.444 hope you can understand @smartstudent

OpenStudy (anonymous):

As a decimal what would that be? @bhaskar.pathk

OpenStudy (anonymous):

it would be 4.44 (round off)

OpenStudy (anonymous):

Im sorry lol I meant a fraction @bhaskar.pathk

OpenStudy (anonymous):

it would be (40/9 ) only